# How Worst-Case Are Adversarial Attacks? Linking Adversarial and Statistical Robustness

### v1

@@ Appendix C Results on CIFAR-10
Following the same evaluation protocol adopted for ImageNet in Section 6, we analyze the directional noisy-risk metric at $\kappa=\kappa^{*}$, denoted by $\bar{\mathcal{R}}_{\kappa^{*}}$, as well as the smallest value of $\kappa$ for which $\bar{\mathcal{R}}=0.25$. The table also includes an explicit ablation over the adversarial concentration parameter $\kappa_{\text{adv}}$, obtained by considering different variants of the DN attack. In this case, $\kappa_{\text{adv}}$ is explored over a lower range compared to Section 6.5, reflecting the reduced input dimensionality of CIFAR-10. As a consequence, the suboptimal or transition values of $\kappa_{\text{adv}}$ differ from those observed in Table 1, consistently with the dimensional scaling discussed in Appendix B.
Overall, the CIFAR-10 results further support the proposed analysis and confirm that the directional noisy-risk metric provides a more informative characterization of attack effectiveness in low-$\kappa$ regimes. In particular, when $\kappa_{\text{adv}}$ is close to $\kappa^{*}$, the metric highlights whether an adversarial perturbation corresponds to a statistically representative noisy failure region, rather than an isolated worst-case direction. Consistently with the ImageNet analysis, we also observe a clear trade-off between adversarial success rate and robustness under stochastic perturbations. While strong iterative attacks such as PGD-5 and PGD-20 achieve high adversarial success rates, they often exhibit limited effectiveness in low-$\kappa$ regimes, indicating that their success is concentrated along highly localized directions. In contrast, attacks that maintain higher noisy-risk at low $\kappa$ correspond to broader and more stable failure regions, which are more relevant from a safety-oriented perspective.
We also report in Figure  9 the analysis of $\bar{\mathcal{R}}$ for attacks with $p=2$ as $\kappa$ varies, focusing on small $\kappa$ values that are closer to the informative noise regime (see Section 6.3). Consistently with the ImageNet analysis, the use of DN attacks exhibits a steeper increase in success rate in these low-$\kappa$ regimes. At the same time, for all attacks, the curves clearly highlight the gap with respect to the classical adversarial success rate observed in the limit $\kappa\to\infty$, as reported in Table 2.
@@ Appendix D Confidence Score Analysis
We further analyze the confidence-score trends introduced in Section 6.4 also for two additional models, WideResNet-101 and VGG-16. Figure 11 reports the corresponding results. Consistent with the results in the paper, we observe a clear relationship between directional noisy risk and model confidence. In particular, adversarial attacks that exhibit high noisy-risk values (i.e., large $\mathcal{R}$) already at low $\kappa$ are strongly associated with lower prediction confidence. This behavior remarks the trend and indicates that directions inducing statistically representative noisy failure regions tend to reduce model confidence more substantially than highly localized, worst-case perturbations, further supporting the proposed interpretation of the directional noisy-risk metric.

### v2

@@ Appendix C Results on CIFAR-10
Following the same evaluation protocol adopted for ImageNet in Section 6, we analyze the directional perturbation risk metric at $\kappa=\kappa^{*}$, denoted by $\bar{\mathcal{R}}_{\kappa^{*}}$, as well as the smallest value of $\kappa$ for which $\bar{\mathcal{R}}=0.25$. The table also includes an explicit ablation over the adversarial concentration parameter $\kappa_{\text{adv}}$, obtained by considering different variants of the DN attack. In this case, $\kappa_{\text{adv}}$ is explored over a lower range compared to Section 6.5, reflecting the reduced input dimensionality of CIFAR-10. As a consequence, the suboptimal or transition values of $\kappa_{\text{adv}}$ differ from those observed in Table 1, consistently with the dimensional scaling discussed in Appendix B.
Overall, the CIFAR-10 results further support the proposed analysis and confirm that the directional perturbation risk provides a more informative characterization of attack effectiveness in low-$\kappa$ regimes. In particular, when $\kappa_{\text{adv}}$ is close to $\kappa^{*}$, the metric highlights whether an adversarial perturbation corresponds to a statistically representative noisy failure region, rather than an isolated worst-case direction. Consistent with the ImageNet analysis, we also observe a clear trade-off between adversarial success rate and robustness under stochastic perturbations. While strong iterative attacks such as PGD-5 and PGD-20 achieve high adversarial success rates, they often exhibit limited effectiveness in low-$\kappa$ regimes, indicating that their success is concentrated along highly localized directions. In contrast, attacks that maintain higher perturbation risk at low $\kappa$ correspond to broader and more stable failure regions, which are more relevant from a safety-oriented perspective.
We also report in Figure  9 the analysis of $\bar{\mathcal{R}}$ for attacks with $p=2$ as $\kappa$ varies, focusing on small $\kappa$ values that are closer to the informative noise regime (see Section 6.3). Consistent with the ImageNet analysis, the use of DN attacks exhibits a steeper increase in success rate in these low-$\kappa$ regimes. At the same time, for all attacks, the curves clearly highlight the gap with respect to the classical adversarial success rate observed in the limit $\kappa\to\infty$, as reported in Table 2.
@@ Appendix D Confidence Score Analysis
We further analyze the confidence-score trends introduced in Section 6.4 also for two additional models, WideResNet-101 and VGG-16. Figure 11 reports the corresponding results. Consistent with the results in the paper, we observe a clear relationship between directional perturbation risk and model confidence. In particular, adversarial attacks that exhibit high perturbation risk values (i.e., large $\mathcal{R}$) already at low $\kappa$ are strongly associated with lower prediction confidence. This behavior remarks the trend and indicates that directions inducing statistically representative noisy failure regions tend to reduce model confidence more substantially than highly localized, worst-case perturbations, further supporting the proposed interpretation of the directional perturbation risk.